The president of Doerman Distributors, Inc., believes that 28% of the firm's orders come from first-time customers. A simple random sample of 100 orders will be used to estimate the proportion of first-time customers. Use z-table.

What is the probability that the sample proportion will be between 0.16 and 0.40 (to 4 decimals)?

What is the probability that the sample proportion will be between 0.21 and 0.35 (to 4 decimals)?
Answer #1

1)

for normal distribution z score =(p̂-p)/σp
here population proportion=     p= 0.280
sample size       =n= 100
std error of proportion=σp=√(p*(1-p)/n)= 0.0449
probability = P(0.16<X<0.4) = P(-2.67<Z<2.67)= 0.9962-0.0038= 0.9924

2)

probability = P(0.21<X<0.35) = P(-1.56<Z<1.56)= 0.9406-0.0594= 0.8812